The surface area of a pyramid with a 7 in. x 7 in. square base is 105 ft sq. What is the slant height?
Answer by ewatrrr(24785)


Hi
The surface area of a pyramid with a 7 in. x 7 in. square base is 105 ft sq.
7*7 = 49 105-49 = 56 = 4(1/2)7s (where s is the slant height of the four triangular faces)
56 = 4(1/2)7s
56/14 = s
s = 4in